(a) What magnitude point charge creates a 10,000 N Please show work and formulas used for credit
Answer #1

(a) The magnitude of electric field of a point charge at of charge q a distance r from it is given by

E=\frac{1}{4\pi\epsilon_o}\frac{q}{r^2}=9\times10^9\frac{q}{r^2}

Since we have to find q for a given E and r, we rearrange the above formula to make q the subject. Hence we get

q=\frac{E\times r^2}{9\times10^9}=\frac{10000\times0.25^2}{9\times10^9}=6.944\times10^{-8}~\mathrm{C}

(b) The electric field at the distance r=10~\mathrm{m} is given by

E=\frac{1}{4\pi\epsilon_o}\frac{q}{r^2}=\frac{9\times10^9\times6.944\times10^{-8}}{10^2}=6.2496~\mathrm{N/C}
Answer #1

(a) The magnitude of electric field of a point charge at of charge q a distance r from it is given by

E=\frac{1}{4\pi\epsilon_o}\frac{q}{r^2}=9\times10^9\frac{q}{r^2}

Since we have to find q for a given E and r, we rearrange the above formula to make q the subject. Hence we get

q=\frac{E\times r^2}{9\times10^9}=\frac{10000\times0.25^2}{9\times10^9}=6.944\times10^{-8}~\mathrm{C}

(b) The electric field at the distance r=10~\mathrm{m} is given by

E=\frac{1}{4\pi\epsilon_o}\frac{q}{r^2}=\frac{9\times10^9\times6.944\times10^{-8}}{10^2}=6.2496~\mathrm{N/C}
Answer #1

(a) The magnitude of electric field of a point charge at of charge q a distance r from it is given by

E=\frac{1}{4\pi\epsilon_o}\frac{q}{r^2}=9\times10^9\frac{q}{r^2}

Since we have to find q for a given E and r, we rearrange the above formula to make q the subject. Hence we get

q=\frac{E\times r^2}{9\times10^9}=\frac{10000\times0.25^2}{9\times10^9}=6.944\times10^{-8}~\mathrm{C}

(b) The electric field at the distance r=10~\mathrm{m} is given by

E=\frac{1}{4\pi\epsilon_o}\frac{q}{r^2}=\frac{9\times10^9\times6.944\times10^{-8}}{10^2}=6.2496~\mathrm{N/C}